At the beginning of 2013, the Harding Construction Company received a contract to build an office building for $10 million. Harding will construct the building according to specifications provided by the buyer, and the project is estimated to take three years to complete. According to the contract, Harding will bill the buyer in installments over the construction period according to a prearranged schedule. Information related to the contract is as follows:

2013

2014

2015

Cost incurred during the year

$2,300,000

$3,600,000

$2,100,000

Estimated costs to complete

5,300,000

2,000,000

0

Billings during the year

1,700,000

4,000,000

4,300,000

Cash collections during the year

1,600,000

3,600,000

4,300,000

  

Calculate the following:

Gross profit recognized:

Percentage of completion Method

2013

2014

2015

Total gross profit:

Please show all the process and correct answers
Answer #1

Answer------

Gross profit recognized: Percentage of completion Method Completed contract Method
2013 $       7,26,316
2014 $       8,42,039
2015 $       4,31,646 $     20,00,000
Total gross profit $     20,00,000 $     20,00,000

Working for percentage of completion

Working 2013 2014 2015
A Contract Price $ 10,000,000 $ 10,000,000 $ 10,000,000
B Cost Incurred to Date $ 2,300,000 $ 5,900,000 $ 8,000,000
C Estimated cost yet to be incurred to complete the contract $ 5,300,000 $ 2,000,000 $ 0
D = B+C Total Cost $ 7,600,000 $ 7,900,000 $ 8,000,000
E = (D/B) x 100 % of Complition 30% 75% 100%
F = A x E Revenue to date $ 3,026,316 $ 7,468,354 $ 10,000,000
G Revenue of Previous year $ 0 $ 3,026,316 $ 7,468,354
H = F - G Net Revenue this year $ 3,026,316 $ 4,442,039 $ 2,531,646
I Cost to date $ 2,300,000 $ 5,900,000 $ 8,000,000
J Cost to date of previous year $ 0 $ 2,300,000 $ 5,900,000
K = I - J Net Cost for the year $ 2,300,000 $ 3,600,000 $ 2,100,000
L = H - K Gross Profits $ 726,316 $ 842,039 $ 431,646

In case of Completed contract method all revenues and gross profit is recognized at the time the contract is complete.